Margarette is twice as old as Isabela. If the age of Margarette is added fo the half the age of Isabela, the result is 60. What are their ages 7 years ago? (2 variables and full solution please, thanks!)
Answer by Theo(13342)

let m = the age of margarette.
let i = the age of isabela.
your equations are:
m = 2i (age of margarette is twice the age of isabela).
m + i/2 = 60 (age of margarette plus half the age of isabela = 60).
in the second equation, replace m with 2i to get:
2i + i/2 = 60
multiply both sides of this equation by 2 to get:
4i + i = 120
combine like terms to get:
5i = 120
solve for i to get:
i = 120/5 = 24.
since m = 2i, then m = 48
you have:
m = 48
i = 24
m = 2 * i becomes 48 = 2 * 24 which becomes 48 = 48 which is true.
m + i/2 = 60 becomes 48 + 24/2 = 60 which becomes 48 + 12 = 60 which becomes 60 = 60 which is true.
numbers check out.
margarette is 48 today.
isabela is 24 today.
7 years ago, margarette was 48 - 7 = 41
7 years ago, isabela was 24 - 7 = 17.
that's your solution.